OpenStudy (anonymous):

Find the inverse of this equation f(x)= 3x + 2 Explain the process to me ^__<

OpenStudy (anonymous):

1) Swap your f(x) and x around so x=3(f(x))+2 2) Rearrange and solve for f(x) That's it

OpenStudy (isaiah.feynman):

To find the inverse of a function Change"f(x)" to "y" Switch x and y Solve for y The new expression is your inverse.

OpenStudy (nincompoop):

generally, another way of saying y is f(x) and vice versa so, put y for f(x) y = 3x + 2 add (-2) on both sides of the equation y+(-2) = 3x divide both sides by 3 (y-2)/3 = x swap sides x = (y-2)/3 put f(y)^-1 for x f(y)^1 = (y-2)/3

OpenStudy (nincompoop):

if you followed Mr. Feynman's way it works the same y = 3x + 2 swap x and y x = 3y + 2 solve for y add (-2) on both sides x+(-2) = 3y divide both sides by 3 (x-2)/3 = y your inverse function is denoted by the notation f^-1, but this time instead of f^-1(y) we will use x, because we swapped x and y earlier. therefore, the inverse of f(x) is f^1(x) = (x-2)/3

OpenStudy (isaiah.feynman):

I would feel honored to be called Mr. Feynman. Lol @nincompoop

OpenStudy (nincompoop):

laughing out loud

OpenStudy (isaiah.feynman):

No you aren't

OpenStudy (nincompoop):

laughing out loud silently ...

OpenStudy (isaiah.feynman):

You do know who Feynman was right?

OpenStudy (nincompoop):

I am a huge fan of Richard P. Feynman btw he authored my first physics book

OpenStudy (isaiah.feynman):

OH MINE! I'M THE ONLY FEYNMAN FAN I KNOW RIGHT UNTIL THIS MOMENT!!!

OpenStudy (nincompoop):

a bit verbose, but it's fine because he's making point that there's no absolute and they aren't tangents (linguistics, not geometrical position)

OpenStudy (isaiah.feynman):

Btw what's the name of your first physics book?